# Vision Foundation Models as Effective Visual Tokenizers for Autoregressive Generation

### v1

@@ Abstract
Leveraging the powerful representations of pre-trained vision foundation models– traditionally used for visual comprehension–we explore a novel direction: building an image tokenizer directly atop such models, a largely underexplored area. Specifically, we employ a frozen vision foundation model as the encoder of our tokenizer. To enhance its effectiveness, we introduce two key components: (1) a region-adaptive quantization framework that reduces redundancy in the pre-trained features on regular 2D grids, and (2) a semantic reconstruction objective that aligns the tokenizer’s outputs with the foundation model’s representations to preserve semantic fidelity. Based on these designs, our proposed image tokenizer, VFMTok, achieves substantial improvements in image reconstruction and generation quality, while also enhancing token efficiency. It further boosts autoregressive (AR) generation—achieving a gFID of 2.07 on ImageNet benchmarks, while accelerating model convergence by three times, and enabling high-fidelity class-conditional synthesis without the need for classifier-free guidance (CFG). The code will be released publicly to benefit the community.
@@ 1 Introduction
GPT’s success in language generation has spurred interest in autoregressive (AR) image generation [44, 43, 50], which relies on visual tokenizers like VQGAN [48, 39, 13, 52, 43] to map images into compact, discrete latent spaces. However, these tokenizers, typically trained from scratch and optimized for reconstruction, often yield latent spaces rich in low-level details but poor in high-level semantics and laden with redundancy. Such flawed latent spaces not only prolong AR model training (Fig. 1) but also necessitate techniques like classifier-free guidance (CFG) for high-fidelity, class-conditional image generation, which in turn increases inference time.
@@
Can VFMs be effective tokenizers? To address this, we initialized the encoder of a VQGAN with different frozen pre-trained foundation models to reconstruct images. Once trained, the tokenizer is integrated on top of an AR model for image synthesis (implementation details depicted in Sec. 3.2) As shown in Tab. 1 (middle rows), our results demonstrate that the semantically rich features from these foundation models not only support effective image reconstruction but also achieve generative performance comparable to—or even surpassing—that of a fully trained VQGAN encoder optimized for both reconstruction and generation. These findings highlight the strong potential of pre-trained vision foundation models to serve as efficient and effective tokenizers for image generation tasks, eliminating the need for extensive encoder training while improving qualities.
Can we improve token efficiency for VFMs? Building on this pilot study, we are further motivated by the observation that natural images often consist of irregular regions that exhibit recurring visual patterns. For example, as illustrated in Fig. 1(a), the upper portion of the crystal ball exhibits consistent patterns such as texture and transparency; similarly, the moss in the stone possesses similar textural structure. When such images are represented using a regular 2D feature grid extracted from foundation models, this structure-agnostic representation may introduce significant redundancy. Exploiting redundancy within semantically coherent regions offers a promising direction for improving tokenization efficiency. Motivated by this insight, we propose a region-adaptive strategy to refine the latent space that aims to enhance both image reconstruction and generation quality while significantly improving token representation efficiency.
@@
Our solution and results. Guided by the preceding experimental analysis and insights, we introduce VFMTok, an image tokenizer that leverages a frozen pre-trained vision foundation model for adaptive region-level tokenization. VFMTok is designed to achieve high reconstruction and generation quality with improved token efficiency. Specifically, VFMTok employs a frozen pre-trained VFM as an encoder to extract multi-level semantic features. A set of learnable anchor queries performs region-level sampling on these features via deformable attention [62], producing region-adaptive tokens that are subsequently quantized into discrete tokens representing the image’s latent representation. These contextual tokens are then processed by a lightweight Vision Transformer [12](ViT) in a BERT-style framework [11, 15] with two primary reconstruction objectives. First, the original image pixels are reconstructed after de-quantization using a VQGAN [43] decoder. Then, the model reconstructs the features from the frozen foundation model itself, allowing VFMTok to retain the semantic richness and discriminative power of the original representations. Once trained, VFMTok enables standard autoregressive Transformers (e.g., Llama [45]) to generate contextual token sequences, which are decoded back into images via the VQGAN decoder, facilitating high-quality image synthesis with compact and semantically meaningful representations. As shown in Tab. 1 (bottom rows), VFMTok achieves superior reconstruction and generation performance while using fewer than half the original number of tokens (256 vs. 576).
@@
Extensive experiments validate that VFMTok, by combining the representational power of visual foundation models with a novel region-adaptive tokenization strategy based on irregular sampling and learnable anchor queries, enables both high-quality and efficient image reconstruction and autoregressive (AR) generation. First, VFMTok achieves superior reconstruction quality and captures richer semantics using significantly fewer tokens compared to prior methods (e.g., 256 vs. 576 in [43]), resulting in a structured, semantic-aware, and compact latent space. As shown in Tab. 1, VFMTok, with only 256 tokens, outperforms other tokenizers using the same VFM encoder by delivering superior reconstruction quality and stronger semantic representation (as indicated by linear probing). Second, the high-quality latent space produced by VFMTok facilitates effective AR training using a simple LLaMA-based model, leading to faster convergence (see Fig. 1(b)) and improved generation performance. Notably, the 1.4B AR model surpasses the performance of LlamaGen-3B despite having fewer parameters and requiring fewer training iterations. The 3B model achieves a new state-of-the-art with a gFID of 2.07 on ImageNet [10] $256\times 256$, outperforming widely-used diffusion models. Third, due to the compact token space and the reduced number of tokens, VFMTok significantly improves the inference speed of AR models (see Tab. 1). Moreover, the rich semantic content embedded in the latent tokens enables effective class-conditional image synthesis with high fidelity—without the need for classifier-free guidance—further reducing inference time.
@@ 3.1 Preliminary
$\begin{split}{Z_{2D}}&=\mathcal{E}(\text{I})\,,\\ \mathcal{VQ}({z_{i}})={c_{i}},\quad\text{where}\quad{i}&=\mathop{\arg\min}\limits_{{j}\in\{1,2,...,N\}}\|{z_{i}}-{c_{j}}\|_{2}\,,\end{split}$ (1)
where $H$ and $W$ denote the input image’s height and width, respectively. $D$ depicts the latent feature dimension. Once discrete tokens are acquired, they can be de-quantized into corresponding code and converted back to image pixels by the decoder $\mathcal{D}(\cdot)$, as depicted in Eq. 2.
@@
To optimize the codebook, the training objective is $\mathbf{\mathcal{L}_{vq}}=\sum{\|\mathbf{sg}(z_{i})-{c_{i}}\|^{2}_{2}+{\beta}\cdot\|\mathbf{sg}({c_{i}})-{z_{i}}\|^{2}_{2}}$, where $\mathbf{sg}(\cdot)$ is a stop-gradient function [2, 48]. The second term is a commitment loss weighted by $\beta$ to align extracted features with codebook vectors. For image reconstruction, the loss function is $\mathcal{L}_{AE}=\mathcal{L}_{2}(\text{I},\hat{\text{I}})+\mathcal{L}_{P}(\text{I},\hat{\text{I}})+\lambda_{G}\cdot\mathcal{L}_{G}(\hat{\text{I}})$, where $\mathcal{L}_{2}$ is a pixel-wise reconstruction loss, $\mathcal{L}_{P}$ is perceptual loss from LPIPS [58], and $\mathcal{L}_{G}$ is adversarial loss from PatchGAN [19] with weight $\lambda_{G}$.
@@ 3.2 Pilot Study: Pre-trained Vision Foundation Models as Tokenizers for AR Generation
As Tab. 1 illustrates, directly using features from pre-trained VFMs yields decent image reconstruction and generation performance compared to vanilla VQGANs. Notably, these VFM-based tokenizers consistently exhibit stronger semantic representation capabilities (as indicated by the linear probing experiment in Tab. 1). For instance, VQGAN (SigLIP2) achieves reconstruction performance on par with vanilla VQGAN, while exhibiting better semantic representation and superior generation quality. Nevertheless, variations in image reconstruction and generation quality arise when different VFMs are used to initialize the tokenizer’s encoder. Specifically, VQGAN (DINOv2) and VQGAN (SigLIP2) demonstrate similar reconstruction and generation quality, both outperforming vanilla VQGAN, while the reconstruction quality of VQGAN (CLIP) trails that of vanilla VQGAN. One contributing factor is that different learning objectives used to train VFMs influence their ability to extract detailed and semantic features from images, thereby affecting downstream image reconstruction and generation quality. As evidence, both DINOv2 [9] and SigLIP2 [46] employed a masked prediction objective to optimize their VFMs, whereas CLIP [37] did not.
@@ 3.3 VFMTok
Decoder of VFMTok for Image and VFM Feature Reconstruction. After de-quantization, the region-adaptive tokens $\tilde{Z}_{r}$ are used for image reconstruction. Since these tokens represent irregular, region-level features, decoding them into a regular 2D image grid requires alignment. To achieve this, we introduce a set of mask tokens $M_{\text{I}}$, representing a 2D feature map of size $H_{m}\times W_{m}$ with channel dimension $C$. The mask tokens are initialized by replicating a single learnable token $H_{m}\times W_{m}$ times. Position embeddings $E$, encoding spatial locations, are then added to form position-aware masked tokens. Next, the de-quantized region-adaptive tokens $\tilde{Z}$ are concatenated with $M_{\text{I}}$, and the combined sequence is processed by a lightweight Transformer $\mathcal{E}_{\text{ViT}}(\cdot)$, which propagates information from the region-adaptive tokens to the masked image tokens. This Transformer employs causal self-attention, aligning its latent space with the structure of autoregressive models. Following DINOv2 [9], we further enrich the input sequence by appending a [CLS]monospace-$[CLS]+$] token and several register tokens to improve representation learning and capture global context—though these are not used for reconstruction. The output of this Transformer is a refined set of mask tokens $\mathcal{F}_{I}$ representing a regular 2D grid structure. These are reshaped into a spatial grid and passed into a decoder $\mathcal{D}(\cdot)$ to reconstruct the image.
@@ 4.1 Setup
Class-conditional Autoregressive Image Generation. Following the generation procedure in LlamaGen [43], the AR models first generate images of $336\times 336$ and then resize them to $256\times 256$ for evaluation. Considering computational costs, we set the training duration based on the number of models’ parameters. Models with fewer than 1B parameters are trained for 300 epochs, while the remaining models are trained for 200 epochs. Beyond the resolution and training duration, all models are trained with the same settings as LlamaGen [43]. Furthermore, in our experiments, AR generation is conducted with both classifier-free guidance (CFG) and a CFG-free protocol.
@@
Table 2: Comparison with other image tokenizers. ${}^{\text{oim.}}$ indicates trained on OpenImages [22]. $\mathcal{Q}_{c}$/$\mathcal{Q}_{P}$ denotes the codebook usage in contextual and patch-level quantizers, respectively.
Method Tokenizer Setup Image Recon. Usage (%$\uparrow$
@@ 4.2 Main Results
Class-conditional Image Generation. We evaluate VFMTok on autoregressive generation models with parameters of  [^1] M (VFMTok-B), 343M(VFMTok-L), 1.4B (VFMTok-XXL), 2.2B (VFMTok-2B) and 3.1B (VFMTok-3B) on ImageNet [10] $256\times 256$ class-conditional image generation task, and compare them with the mainstream generation models, including diffusion models (Diff.) [60, 35, 31, 51], BERT-style masked-prediction models (Mask.) [4], and AR generation models (AR) [55, 13, 52, 24, 39, 43].
As shown in Tab. 3, our models exhibit competitive performance across all metrics compared to mainstream image generation models. Notably, VFMTok beats BERT-style models [4] in terms of gFID without the requirement of complicated sampling tuning. With comparable or even fewer parameters, our method surpasses most AR generative models [55, 13, 52, 24, 39] in both gFID and gIS metrics. Under the same training setting, VFMTok surpasses LlamaGen [43] by significant gFID gains and notable gIS improvements. Specifically, VFMTok-B outperforms LlamaGen-B [43] with gains of 2.56 in gFID and 69.7 in gIS. Besides, our VFMTok-L model achieves a gFID of 2.75 at 300 epochs, also obtaining a gain of 22.7 in gIS. Notably, when compared with LlamaGen-3B with 3B parameters, our VFMTok-XXL achieves even better generation performance with less than half the number of parameters and fewer training iterations. Additionally, class-conditional image generation results are visualized in the Appendix.
@@ 4.3 Ablation Study and More Analysis
Component study. To assess the contribution of each proposed component to image reconstruction and synthesis, we conduct a step-by-step component analysis using a baseline tokenizer built on vanilla VQGAN [43]. We incrementally add the following components: (1) replace the VQGAN encoder with a frozen pre-trained foundation model (DINOv2-L [9]); (2) introduce learnable queries and a deformable attention for region-adaptive tokenization, using only single-level features from the final layer; (3) incorporate multi-level features to enrich representations with both low-level detail and high-level semantics; and (4) add a feature reconstruction objective based on pre-trained VFM outputs. After training each tokenizer, we integrate it with our AR generation model, VFMTok-L, for autoregressive image synthesis. Both the tokenizer and AR model are trained for 50 epochs. Additionally, we perform linear probing on the [CLS]monospace-$[CLS]+$] token, following the MAE [15] protocol.
As shown in Tab. 4, replacing VQGAN’s encoder with a frozen pre-trained vision foundation model yields reconstruction and generation performance on par with a VQGAN trained specifically for visual reconstruction using 576 tokens. This substitution also significantly enhances the semantic quality of the tokenizer’s representations. To further improve token efficiency, we introduce region-adaptive tokenization using deformable attention to exploit the spatial redundancy inherent in regular 2D grid features. This reduces the number of visual tokens to 256. However, this performance gain comes at a cost: reconstruction and generation quality degrade slightly due to two factors: (1) fewer visual tokens limit representational capacity, and (2) the absence of explicit supervision hinders the effective optimization of the region-adaptive tokens. To address this, we incorporate multi-level feature extraction, which improves the reconstruction capability by leveraging both low- and high-level information. However, without additional guidance, the semantic consistency of the learned tokens may still degrade. Finally, we introduce a pre-trained feature reconstruction objective, which significantly boosts both image reconstruction and generation quality. This objective encourages alignment with semantic features from the frozen VFM and effectively balances the contributions of low- and high-level features to the contextual tokens—thereby preserving semantic fidelity.
@@
Convergence and efficiency analysis. Beyond above analysis, we experiment VFMTok with a randomly initialized encoder instead of a pre-trained VFM with other components remaining unchanged. As shown in Tab. 4 (last row), its reconstruction quality dropped to the level of VQGAN. Meanwhile, both its semantic representation capability and generation performance also decreased. This indicates a frozen VFM benefits tokenizer training as it provides a latent space advantageous for image reconstruction and generation. Besides, those semantic-rich, structured latent tokens accelerate AR model training convergence. As evidenced in Fig. 1(b), VFMTok enables AR models to achieve a 3$\times$ speedup in convergence compared to VQGAN. Moreover, an AR model’s generation time is quadratically proportional to the number of tokens. At the same resolution, VFMTok uses approximately half the tokens for image representation compared to counterparts like DINOv2-VQGAN and CLIP-VQGAN. Consequently, VFMTok achieves a 4$\times$ generation speedup over these counterparts depicted in Tab. 1. This acceleration can be further enhanced with CFG-free generation.

### v2

@@ Abstract
In this work, we present a novel direction to build an image tokenizer directly on top of a frozen vision foundation model, which is a largely underexplored area. Specifically, we employ a frozen vision foundation model as the encoder of our tokenizer. To enhance its effectiveness, we introduce two key components: (1) a region-adaptive quantization framework that reduces redundancy in the pre-trained features on regular 2D grids, and (2) a semantic reconstruction objective that aligns the tokenizer’s outputs with the foundation model’s representations to preserve semantic fidelity. Based on these designs, our proposed image tokenizer, VFMTok, achieves substantial improvements in image reconstruction and generation quality, while also enhancing token efficiency. It further boosts autoregressive (AR) generation—achieving a gFID of 1.36 on ImageNet benchmarks, while accelerating model convergence by three times, and enabling high-fidelity class-conditional synthesis without the need for classifier-free guidance (CFG). The code is available at https://github.com/CVMI-Lab/VFMTok.
@@ 1 Introduction
GPT’s success in language generation has spurred interest in autoregressive (AR) image generation [44, 43, 50], which relies on visual tokenizers like VQGAN [48, 39, 13, 52, 43] to map images into compact, discrete latent spaces. However, these tokenizers, typically trained from scratch and optimized for reconstruction, often yield latent spaces rich in low-level details but poor in high-level semantics and laden with redundancy. Such flawed latent spaces not only prolong AR model training (Fig.˜1) but also necessitate techniques like classifier-free guidance (CFG) for high-fidelity, class-conditional image generation, which in turn increases inference time.
@@
Can VFMs be effective tokenizers? To address this, we initialized the encoder of a VQGAN with different frozen pre-trained foundation models to reconstruct images. Once trained, the tokenizer is integrated on top of an AR model for image synthesis (implementation details depicted in Sec.˜3.2) As shown in Tab.˜1 (middle rows), our results demonstrate that the semantically rich features from these foundation models not only support effective image reconstruction but also achieve generative performance comparable to—or even surpassing—that of a fully trained VQGAN encoder optimized for both reconstruction and generation. These findings highlight the strong potential of pre-trained vision foundation models to serve as efficient and effective tokenizers for image generation tasks, eliminating the need for extensive encoder training while improving qualities.
Can we improve token efficiency for VFMs? Building on this pilot study, we are further motivated by the observation that natural images often consist of irregular regions that exhibit recurring visual patterns. For example, as illustrated in Fig.˜1(a), the upper portion of the crystal ball exhibits consistent patterns such as texture and transparency; similarly, the moss in the stone possesses similar textural structure. When such images are represented using a regular 2D feature grid extracted from foundation models, this structure-agnostic representation may introduce significant redundancy. Exploiting redundancy within semantically coherent regions offers a promising direction for improving tokenization efficiency. Motivated by this insight, we propose a region-adaptive strategy to refine the latent space that aims to enhance both image reconstruction and generation quality while significantly improving token representation efficiency.
@@
Our solution and results. Guided by the preceding experimental analysis and insights, we introduce VFMTok, an image tokenizer that leverages a frozen pre-trained vision foundation model for adaptive region-level tokenization. VFMTok is designed to achieve high reconstruction and generation quality with improved token efficiency. Specifically, VFMTok employs a frozen pre-trained VFM as an encoder to extract multi-level semantic features. A set of learnable anchor queries performs region-level sampling on these features via deformable attention [64], producing region-adaptive tokens that are subsequently quantized into discrete tokens representing the image’s latent representation. These contextual tokens are then processed by a lightweight Vision Transformer [12](ViT) in a BERT-style framework [11, 15] with two primary reconstruction objectives. First, the original image pixels are reconstructed after de-quantization using a VQGAN [43] decoder. Then, the model reconstructs the features from the frozen foundation model itself, allowing VFMTok to retain the semantic richness and discriminative power of the original representations. Once trained, VFMTok enables standard autoregressive Transformers (e.g., Llama [45]) to generate contextual token sequences, which are decoded back into images via the VQGAN decoder, facilitating high-quality image synthesis with compact and semantically meaningful representations. As shown in Tab.˜1 (bottom rows), VFMTok achieves superior reconstruction and generation performance while using fewer than half the original number of tokens (256 vs. 576).
@@
Extensive experiments validate that VFMTok, by combining the representational power of visual foundation models with a novel region-adaptive tokenization strategy based on irregular sampling and learnable anchor queries, enables both high-quality and efficient image reconstruction and autoregressive (AR) generation. First, VFMTok achieves superior reconstruction quality and captures richer semantics using significantly fewer tokens compared to prior methods (e.g., 256 vs. 576 in [43]), resulting in a structured, semantic-aware, and compact latent space. As shown in Tab.˜1, VFMTok, with only 256 tokens, outperforms other tokenizers using the same VFM encoder by delivering superior reconstruction quality and stronger semantic representation (as indicated by linear probing). Second, the high-quality latent space produced by VFMTok facilitates effective AR training using a simple LLaMA-based model, leading to faster convergence (see Fig.˜1(b)) and improved generation performance. Notably, the 1.4B AR model surpasses the performance of LlamaGen-3B despite having fewer parameters and requiring fewer training iterations. The 1.5B advanced AR model achieves a new state-of-the-art with a gFID of 1.36 on ImageNet [10] $256\times 256$, outperforming widely-used diffusion models. Third, due to the compact token space and the reduced number of tokens, VFMTok significantly improves the inference speed of AR models (see Tab.˜1). Moreover, the rich semantic content embedded in the latent tokens enables effective class-conditional image synthesis with high fidelity—without the need for classifier-free guidance—further reducing inference time.
@@ 3.1 Preliminary
[…]
where $H$ and $W$ denote the input image’s height and width, respectively. $D$ depicts the latent feature dimension. Once discrete tokens are acquired, they can be de-quantized into corresponding code and converted back to image pixels by the decoder $\mathcal{D}(\cdot)$, as depicted in Eq.˜2.
@@
[…]
@@ 3.2 Pilot Study: Pre-trained Vision Foundation Models as Tokenizers for AR Generation
As Tab.˜1 illustrates, directly using features from pre-trained VFMs yields decent image reconstruction and generation performance compared to vanilla VQGANs. Notably, these VFM-based tokenizers consistently exhibit stronger semantic representation capabilities (as indicated by the linear probing experiment in Tab.˜1). For instance, VQGAN (SigLIP2) achieves reconstruction performance on par with vanilla VQGAN, while exhibiting better semantic representation and superior generation quality. Nevertheless, variations in image reconstruction and generation quality arise when different VFMs are used to initialize the tokenizer’s encoder. Specifically, VQGAN (DINOv2) and VQGAN (SigLIP2) demonstrate similar reconstruction and generation quality, both outperforming vanilla VQGAN, while the reconstruction quality of VQGAN (CLIP) trails that of vanilla VQGAN. One contributing factor is that different learning objectives used to train VFMs influence their ability to extract detailed and semantic features from images, thereby affecting downstream image reconstruction and generation quality. As evidence, both DINOv2 [9] and SigLIP2 [46] employed a masked prediction objective to optimize their VFMs, whereas CLIP [37] did not.
@@ 3.3 VFMTok
Decoder of VFMTok for Image and VFM Feature Reconstruction. After de-quantization, the region-adaptive tokens $\tilde{Z}_{r}$ are used for image reconstruction. Since these tokens represent irregular, region-level features, decoding them into a regular 2D image grid requires alignment. To achieve this, we introduce a set of mask tokens $M_{\text{I}}$, representing a 2D feature map of size $H_{m}\times W_{m}$ with channel dimension $C$. The mask tokens are initialized by replicating a single learnable token $H_{m}\times W_{m}$ times. Position embeddings $E$, encoding spatial locations, are then added to form position-aware masked tokens. Next, the de-quantized region-adaptive tokens $\tilde{Z}$ are concatenated with $M_{\text{I}}$, and the combined sequence is processed by a lightweight Transformer $\mathcal{E}_{\text{ViT}}(\cdot)$, which propagates information from the region-adaptive tokens to the masked image tokens. This Transformer employs causal self-attention, aligning its latent space with the structure of autoregressive models. Following DINOv2 [9], we further enrich the input sequence by appending a $\verb+[CLS]$+ token and several register tokens to improve representation learning and capture global context—though these are not used for reconstruction. The output of this Transformer is a refined set of mask tokens $\mathcal{F}_{I}$ representing a regular 2D grid structure. These are reshaped into a spatial grid and passed into a decoder $\mathcal{D}(\cdot)$ to reconstruct the image.
@@ 4.1 Setup
Class-conditional Autoregressive Image Generation. Following the generation procedure in LlamaGen [43], the AR models first generate images of $336\times 336$ and then resize them to $256\times 256$ for evaluation. Considering computational costs, we set the training duration based on the number of models’ parameters. Models with fewer than 1B parameters are trained for 300 epochs, while the remaining models are trained for 200 epochs. Beyond the resolution and training duration, all models are trained with the same settings as LlamaGen [43]. Furthermore, we also incorporated the same VFMTok into the RAR [55] autoregressive generation framework, with all training settings remaining consistent with RAR [55]. Additionally, in our experiments, AR generation is conducted with both classifier-free guidance (CFG) and a CFG-free protocol.
@@
Table 2: Comparison with other image tokenizers. $^{\text{oim.}}$ indicates trained on OpenImages [22]. $\mathcal{Q}_{c}$$\mathcal{Q}_{P}$ denotes the codebook usage in contextual and patch-level quantizers, respectively.
[…]
@@ 4.2 Main Results
Class-conditional Image Generation. We evaluate VFMTok on vanilla autoregressive models – LlamaGen [43], and advanced generative model – RAR [55] with different scales by conducting $256\times 256$ class-conditional image generation task on ImageNet [10], where comparing them with the mainstream generation models, including diffusion models (Diff.) [61, 35, 31, 51], BERT-style masked-prediction models (Mask.) [4], and AR generation models (AR) [56, 13, 52, 24, 39, 43].
As shown in Tab.˜3, our models exhibit competitive performance across all metrics compared to mainstream image generation models. Notably, VFMTok beats BERT-style models [4] in terms of gFID without the requirement of complicated sampling tuning. With comparable or even fewer parameters, our method surpasses most AR generative models [56, 13, 52, 24, 39] in both gFID and gIS metrics. Under the same training setting, VFMTok surpasses LlamaGen [43] by significant gFID gains and notable gIS improvements. Specifically, VFMTok-B outperforms LlamaGen-B [43] with gains of 2.56 in gFID and 69.7 in gIS. Besides, our VFMTok-L model achieves a gFID of 2.75 at 300 epochs, also obtaining a gain of 22.7 in gIS. Notably, when compared with LlamaGen-3B with 3B parameters, our VFMTok-XXL achieves even better generation performance with less than half the number of parameters and fewer training iterations. Futhermore, when VFMTok is incorporated into RAR [55], it achieves a generative performance with gFID of 1.36, which is the state-of-the-art generation performance at present. Additionally, class-conditional image generation results are visualized in the Appendix.
@@ 4.3 Ablation Study and More Analysis
Component study. To assess the contribution of each proposed component to image reconstruction and synthesis, we conduct a step-by-step component analysis using a baseline tokenizer built on vanilla VQGAN [43]. We incrementally add the following components: (1) replace the VQGAN encoder with a frozen pre-trained foundation model (DINOv2-L [9]); (2) introduce learnable queries and a deformable attention for region-adaptive tokenization, using only single-level features from the final layer; (3) incorporate multi-level features to enrich representations with both low-level detail and high-level semantics; and (4) add a feature reconstruction objective based on pre-trained VFM outputs. After training each tokenizer, we integrate it with our AR generation model, VFMTok-L, for autoregressive image synthesis. Both the tokenizer and AR model are trained for 50 epochs. Additionally, we perform linear probing on the $\verb+[CLS]$+ token, following the MAE [15] protocol.
As shown in Tab.˜4, replacing VQGAN’s encoder with a frozen pre-trained vision foundation model yields reconstruction and generation performance on par with a VQGAN trained specifically for visual reconstruction using 576 tokens. This substitution also significantly enhances the semantic quality of the tokenizer’s representations. To further improve token efficiency, we introduce region-adaptive tokenization using deformable attention to exploit the spatial redundancy inherent in regular 2D grid features. This reduces the number of visual tokens to 256. However, this performance gain comes at a cost: reconstruction and generation quality degrade slightly due to two factors: (1) fewer visual tokens limit representational capacity, and (2) the absence of explicit supervision hinders the effective optimization of the region-adaptive tokens. To address this, we incorporate multi-level feature extraction, which improves the reconstruction capability by leveraging both low- and high-level information. However, without additional guidance, the semantic consistency of the learned tokens may still degrade. Finally, we introduce a pre-trained feature reconstruction objective, which significantly boosts both image reconstruction and generation quality. This objective encourages alignment with semantic features from the frozen VFM and effectively balances the contributions of low- and high-level features to the contextual tokens—thereby preserving semantic fidelity.
@@
[…]